Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

orps1

Expert replies
by stubbornp » Sat Nov 01, 2008 10:42 am
orps1
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by cramya » Sat Nov 01, 2008 11:08 am
Prob is not my forte but is it 1/3


P(NOT BURNING IN FIRST 6 MONTHS) = 1/3

P(BURNING IN FIRST 6 MONTHS) = 1 - 1/3 = 2/3


P(BURNING 6 MONTHS - 1 YR) = 1/2*2/3 = 1/3
Join the discussion

by stop@800 » Sat Nov 01, 2008 11:12 am
first 6 months
burn 1/3
not burn 2/3

next 6 months
not burn = 1/2 prev not burn
= 1/2 * 2/3
= 1/3

so burn = 1-1/3
=2/3
Join the discussion

by cramya » Sat Nov 01, 2008 11:12 am
Juts realized my mistake in the post

P( BURNING IN FIRST 6 MONTHS) = 1/3

P( not BURNING IN FIRST 6 MONTHS) = 1 - 1/3 = 2/3


P(BURNING 6 MONTHS - 1 YR) = 1/2*2/3 = 1/3
p(not burning 6 MONTHS - 1 YR)) = 1-1/3 = 2/3


2/3 E) is my final vote(confirming the fact that prob is not my forte :D )

OA please?
Join the discussion

by cramya » Sat Nov 01, 2008 11:17 am
The first time I read it incorrectlty thinking the given data 1/3 was for not burning.

CARELESS ABD LETHARGIC reading can cost u heavily on GMAT: -) A lesson learnt.....

Good luck All!!!!
Join the discussion

by stubbornp » Sat Nov 01, 2008 1:34 pm
oa is not 2/3
Join the discussion

by stop@800 » Sat Nov 01, 2008 2:50 pm
stubbornp wrote:oa is not 2/3
it shud be 4/9.
Join the discussion

by stubbornp » Sun Nov 02, 2008 9:51 am
stop@800 wrote:
stubbornp wrote:oa is not 2/3
it shud be 4/9.
yup it is 4/9....Please explain
Join the discussion

by cramya » Sun Nov 02, 2008 12:14 pm
Ok Shooting in the dark with these prob problems continues.Stp@800 let me know if this is how u came to the soln 4/9


P(BURNING IN 6 MONTHS-1YR) =
P(NOT BURNING IN THE FIRST 6 MONTHS) * P(BURNING IN 6MONTHS-1YR)

P(NOT BURNING IN FIRST 6 MONTHS) = 1-1/3 = 2/3

P(NOT BURNING IN 6 MONTHS-1YR) = 1-2/3 = 1/3 (WHICH IS 1/2 *1/3 AS GIVEN IN THE PROBLEM)

P(BURNING IN 6 MONTHS-1YR) = 1-1/3 = 2/3

P(BURNING IN 6 MONTHS-1YR) = 2/3*2/3 = 4/9

Is this conditional probanility P(A) * P(B/A) WHERE P(A) is not burning in forst 6 months and P(b/a) is probab of burning in 6 months-1yr given the bulb has not got burnt in first 6 months


Suggestions welcome!!!
Join the discussion

by cramya » Sun Nov 02, 2008 12:18 pm
CORRECTION IN BOLD, Sorry guys I made a meal out of this problem:


Ok Shooting in the dark with these prob problems continues.Stop@800 let me know if this is how u came to the soln 4/9


P(BURNING IN 6 MONTHS-1YR) =
P(NOT BURNING IN THE FIRST 6 MONTHS) * P(BURNING IN 6MONTHS-1YR/NOT BURNING IN THE FIRST 6 MONTHS)

P(NOT BURNING IN FIRST 6 MONTHS) = 1-1/3 = 2/3

P(NOT BURNING IN 6 MONTHS-1YR) = 1/2 *2/3 = 1/3

P(BURNING IN 6 MONTHS-1YR/NOT BURNING IN THE FIRST 6 MONTHS) = 1-1/3 = 2/3


P(BURNING IN 6 MONTHS-1YR) = 2/3*2/3 = 4/9

Is this conditional probability P(A) * P(B/A) WHERE P(A) is not burning in forst 6 months and P(b/a) is probab of burning in 6 months-1yr given the bulb has not got burnt in first 6 months


Suggestions welcome!!!
Join the discussion

by luisengard » Mon Nov 03, 2008 11:10 am
cramya wrote: P(BURNING IN 6 MONTHS-1YR) =
P(NOT BURNING IN THE FIRST 6 MONTHS) * P(BURNING IN 6MONTHS-1YR/NOT BURNING IN THE FIRST 6 MONTHS)
Hi guys,

I could use some help, I'm stuck with probability, permutations and combinations.

I don't understant why you state that the answer is that if the question of the problem is exactly "P(BURNING IN 6MONTHS-1YR/NOT BURNING IN THE FIRST 6 MONTHS)", isn't it?

Thanks a lot!!

(BTW is my first post, thank you very much for the help and this lovely forum!)
Join the discussion

by cramya » Mon Nov 03, 2008 2:38 pm
I could be way off here but hope I am not.

I think since these are dependednt events i.e. P(NOT BURNING IN THE FIRST 6 MONTHS) * P(BURNING IN 6MONTHS-1YR/NOT BURNING IN THE FIRST 6 MONTHS)


To find the prob that abulb will burn in 6 months-1yr we have to take in to account the prob that it hasnot burt in the first 6 months. If it did then this probability would be 0

I am sure prob experts can jump in on luisengard's question. Thanks!
Join the discussion